# LongGenbench: Benchmarking Long-Form Generation in long-context LLMs

### v3

@@ 3.2 Main Result
Significant disparities in performance across models primarily stem from differences in architecture and training datasets. Notably, models with varying parameter sizes, such as Llama3.1-8B-instruction (Dubey et al., 2024) (under 10 billion parameters), Qwen-72B (Yang et al., 2024) (over 20 billion parameters), and GPT-4o-mini (OpenAI, 2024a) (a closed-source model), have demonstrated superior efficacy, successfully completing most primary tasks in full. In contrast, some models struggle with these tasks, exhibiting limitations such as: 1) models responding solely to specified subtasks, neglecting others, and 2) models halting after only completing the initial task segment, despite prompts requiring full sequential subtask completion. This issue may originate from the current instructional tuning data, which could cause partial responses in complex, lengthy tasks. Especially in GPT-4o(OpenAI, 2023), it recognizes that this task will generate a long output and only provides a few examples.
STIC-1 & STIC-2:
The STIC-1 metric demonstrates the robust performance of both the Mixtral-8*7B (Mistral.AI, 2023) and GPT-4o (OpenAI, 2024a) models in adhering to the Scenario Task Instruction (STI) during sub-scenario queries. However, an analysis of the overall STI completion rates reveals subtle performance discrepancies. Models with lower completion rates experience a marked decline from STIC-1 to STIC-2. In this metric, the Mixture of Experts (MOE) models outperformed their dense counterparts.
Length (Number of words):
@@
We calculate the average output word count for models that consistently complete all subtasks, achieving at least an 80% completion rate in sub-scenarios, excluding data from unsuccessful attempts. Most models substantially exceed previous benchmarks for long-form generation tasks in terms of output length. Notably, the LongWriter (Bai et al., 2024) model excels, efficiently meeting word count requirements for each subtask.
Given the results and the weighted average score (wAvg) at a sequence length of 16K, the open-source Qwen2-72B and the closed-source GPT-4o models demonstrate optimal performance. At a sequence length of 32K, the LLama3.1-8B model, known for its excellent completion rate, outperforms models with larger parameters, highlighting its efficiency in managing extended lengths.
@@ 3.3 Accuracy Trend with Varying Sequence Length
As show in Fig. 3, there is a noticeable decline in model performance over time. As the outputs become longer, all models gradually fail to adhere to the initial instructions, resulting in a performance degradation analogous to trends observed in the NIAH dataset. This deviation is particularly pronounced when the output exceeds 4,000 tokens, where adherence to instructions significantly diminishes, with marked deterioration at the 4,000-token threshold and substantial ineffectiveness by 16,000 tokens. Conversely, in tasks such as those involving NIAH and more complex multi-needle tasks of similar lengths, these models demonstrate near-perfect performance. Such disparities in performance cannot be easily explained by prevailing theories, such as ”loss in the middle” (Liu et al., 2024a), especially since the instructions are positioned within the first 500 tokens.
@@
Figure 3: The right side of the figure illustrates the model’s performance on STI tasks at 16K as sequence length increases, whereas the left side depicts performance at 32K. All curves have been smoothed with a Moving Average.
3.4 Three Specific Task Instruction
In Figure 4, we present the performance metrics across various tasks. The model shows similar, yet not identical, proficiency in single and range tasks, suggesting its effective execution of direct instructions for these task types. This reduction in performance can be ascribed to the additional cognitive load associated with interpreting periodic instructions, such as identifying relevant weeks in statements like ”every four weeks starting from week 10.” Consequently, outcomes for periodic tasks are substantially poorer compared to those for single and range tasks, which have well-defined parameters. Typically, the model’s performance hierarchy is represented as $single>range>periodic$.
3.5 Compare with long context input
@@
In this subsection, we examine the relationship between a model’s ability to handle long-range inputs and its performance on long-range outputs. Specifically, we investigate whether a model’s capacity to manage long-range inputs corresponds to improved performance on long-range outputs. For this analysis, we use the RULER dataset. RULER is a novel synthetic benchmark designed to flexibly configure sequence length and task complexity, making it ideal for comprehensive evaluations of long-context language models. We compare the models’ performance on sequences of the same length, as shown in Figure 4, which indicates a certain degree of correlation between input handling and output performance. At 16K, the Pearson correlation coefficient is 0.51, while at 32K, it increases to 0.66, highlighting a significant overlap in their capabilities.
(a) Performance Comparison of Three STI Type
@@
Figure 4: The left fig displays the performance of three different STI models, with the red line representing the average for each category. The right fig shows the performance and correlation of the ruler and LongGenBench  at the same length settings.
4 analysis and limitaion
Richness of Content
Even when sub-scenarios are specifically designed to enhance task richness, the model’s outputs tend to converge as output volume increases. This results in a homogeneity of recorded events despite differences in time and location. Such convergence not only degrades performance, but also diminishes output diversity. Attempts to mitigate this issue by adjusting the parameter $repetition\_penalty$ during model inference have not been effective in our experiments.
Rationality of Content
Our research presently focuses only on evaluating instruction-following capabilities within tasks; a more comprehensive analysis of task rationality and integrity could be a direction for future research. For instance, if a virtual task pertains to a specific profession, all job-related activities recorded in a diary should correspond to that profession. Additionally, we noted that temperature records in diaries fail to reflect realistic temporal changes. For example, in San Francisco, one would expect temperatures to rise from cold (0-10 degrees Celsius) at the beginning of the year to warmer (20-30 degrees Celsius) by mid-year; however, our findings indicate temperatures remain in the warmer range (20-30 degrees Celsius) even into December.
Instruction Data
Significant discrepancies in the performance of most models compared to their ability to process inputs, such as Ruler(Hsieh et al., 2024), are likely due to the length distribution of instruction tuning data. These data sets are typically brief, usually under 200 tokens, and lack extended instructional content. Organizing or synthesizing instruction tuning data with longer outputs could be a potential solution for future research endeavors.
@@ 5 Related Work
Our work is closely related to other works on benchmarking long-context language models. ZeroSCROLLS (Shaham et al., 2023) covers ten realistic natural language tasks, such as long-document QA and (query-based) summarization. L-Eval (An et al., 2024a) also uses realistic data, which was filtered manually to ensure quality. LongBench (Bai et al., 2023b) contains tasks in a bilingual setting. InfiniteBench (Zhang et al., 2024) includes tasks with length greater than 100K tokens. LTM (Castillo et al., 2024) targets the evaluation of long-term conversations. To isolate the effect of parametric knowledge, previous works (Dong et al., 2023; Li et al., 2023b) also propose to use documents posted online later than a certain cutoff date, or leverage extremely low-resource materials (Tanzer et al., 2024). Compared to realistic benchmarks, synthetic tasks are more flexible to control the setup (e.g., sequence length and task complexity) and less affected by parametric knowledge. Recent works have mostly focused on retrieval-based synthetic tasks(Kamradt, 2023; Mohtashami & Jaggi, 2023; Wu et al., 2023; Li et al., 2023a; Liu et al., 2024a), with a few on other types of long-context usage, including various types of reasoning (Tay et al., 2021) and long-range discourse modeling (Sun et al., 2022).
@@
Significant advances have been made in long-form text generation, spanning various applications such as story generation (Fan et al., 2019b; Xu et al., 2020), paragraph completion (Kang & Hovy, 2020), sustained conversation (Xu et al., 2022), comprehensive question answering (Fan et al., 2019a; Dasigi et al., 2021; Stelmakh et al., 2022; Lee et al., 2023), and article production (Hua & Wang, 2020; Hu et al., 2022). Employing large language models (LLMs) for evaluative tasks (Liu et al., 2023; Chiang & Lee, 2023; Liu et al., 2024b; Bai et al., 2024) presents challenges, including limited access to current information and inconsistent performance (Shen et al., 2023). Evaluators often struggle to accurately assess quality in lengthy texts, leading to substantial costs (Gillick & Liu, 2010; Iskender et al., 2020). In contrast, the methodology proposed by Tan et al. (2024) aims to improve the assessment of informativeness and coverage, thus enhancing the evaluation of long-form text generation over extended narratives. However, the average length of text generated in the studies mentioned above does not reach the current context length limitations of long context language models, with most texts being fewer than 2000 tokens.
Instruction following.
Recent advancements in open-ended instruction tuning algorithms and models (Ouyang et al., 2022; Rafailov et al., 2024; OpenAI, 2022; Taori et al., 2023; Chiang et al., 2023) have underscored the need for scalable and cost-effective evaluation methods. Consequently, LLMs have been deployed as evaluators, showing better correlation with human assessments than traditional metrics such as BLEU (Papineni et al., 2002). Despite their advantages, LLMs exhibit biases including sensitivity to the order of presentation and a preference for verbose outputs from models with similar architectures (Wang et al., 2024; Pezeshkpour & Hruschka, 2023; Zheng et al., 2023). This has led to the development of meta-evaluation benchmarks like FairEval, MT-Bench, and LLMEval2(Wang et al., 2024; Zheng et al., 2023; Zhang et al., 2023) to address these biases. Recent studies have focused on improving LLM evaluations by employing diverse evaluators and enhanced prompting strategies(Zheng et al., 2023; Li et al., 2023c; Zhang et al., 2023; Chan et al., 2023). However, it is noteworthy that the length of instruction following in these studies often fails to meet the context length capabilities of current long-context language models.
@@ 6 Conclusion
We present LongGenBench , a synthetic benchmark designed to evaluate the long-form generation capabilities of language models. LongGenBench employs tasks that require following instructions over long distances, offering a comprehensive assessment of language models’ abilities. We evaluated nine advanced language models using LongGenBench, with context sizes ranging from 32K to 128K. Although these models achieved perfect scores in the commonly used ”Ruler”, their performance decreased significantly on LongGenBench. Common failure modes observed include premature termination of the main task, incomplete responses, disregard for prior instructions as outputs lengthened, and repetitive content generation. Our findings indicate that LongGenBench poses significant challenges for even the top-tier open-source models.
Appendix A Models
Our results in the main text only include aligned models (two closed-source model gpt-4o, gpt-4o-mini and eight open-source models, the model information is shown in the table 3.
@@ Appendix A Models
FILM-7B (An et al., 2024b) 7B 128K In2Training/FILM-7B
@@
Appendix B FOUR DISTINCT SCENARIO SETUPS
Here are four templates for different scenarios. In Diary template, the professions and names are interchangeable options.
Appendix C evaluation prompts
evaluation prompts Example 1 Example 2 Context: ”Corresponds to the output part of Specific Task Instruction” Instructions: Does this context include Specific Task Instruction? Please answer with ’yes’ or ’no’ only.”

### v6

@@ Results.
Main Task Completion. Significant disparities in performance across models primarily stem from differences in architecture and training datasets. Notably, models with varying parameter sizes, such as Llama3.1-8B-instruction (Dubey et al., 2024) (under 10 billion parameters), Qwen-72B (Yang et al., 2024) (over 20 billion parameters), and GPT-4o-mini (OpenAI, 2024a) (a closed-source model), have demonstrated superior efficacy, successfully completing most primary tasks in full. In contrast, some models struggle with these tasks, exhibiting limitations such as: 1) models responding solely to specified subtasks, neglecting others, and 2) models halting after only completing the initial task segment, despite prompts requiring full sequential subtask completion. This issue may originate from the current instructional tuning data, which could cause partial responses in complex, lengthy tasks. Especially in GPT-4o(OpenAI, 2023), it recognizes that this task will generate a long output and only provides a few examples.
STIC-1 and STIC-2. The STIC-1 metric revealed strong performance in adhering to task instructions for models like Mixtral-8x7B and GPT-4o-mini, particularly in shorter sequences. However, a significant drop in STIC-2 scores for several models indicates that maintaining instruction adherence over longer text sequences remains a challenge. This performance degradation emphasizes the need for better tuning and architectural modifications to improve long-term coherence. The MoE model, Mistral-8x7B generally outperformed dense counterparts in instruction-following over extended sequences, but both model types struggled with STIC-2 in longer generations.
A common failure mode observed across multiple models was the tendency to forget or misinterpret instructions as the sequence length increased. For example, in the Skyscraper Design task, some models correctly described the initial few floors but deviated from the original plan as the task progressed, particularly in the 32K token setting. This highlights the memory retention issue in long-context models, which often leads to a loss of coherence and adherence to task instructions. Examples of these failures, where models struggled to follow prompt instructions over extended outputs, are provided in Appendix D.
Length (Number of words).
@@
We calculated the average output word count for models that consistently completed all subtasks, achieving at least an 80% completion rate in sub-scenarios, excluding data from unsuccessful attempts. Most models substantially exceeded previous benchmarks for long-form generation tasks in terms of output length. Notably, the LongWriter (Bai et al., 2024) model excelled, efficiently meeting word count requirements for each subtask. Given the results and the weighted average score (wAvg) at a sequence length of 16K, the open-source Qwen2-72B and the closed-source GPT-4o models demonstrated optimal performance. At a sequence length of 32K, the Llama3.1-8B model, outperformed models with larger parameters, highlighting its efficiency in managing extended lengths.
@@ 3.3 Accuracy Trend with Varying Sequence Length
As illustrated in Figure 2, there is a clear decline in model performance as output length increases. Models exhibited strong adherence to initial instructions at shorter sequence lengths, but performance gradually degraded as the text generation extended beyond the 4,000-token threshold. This degradation aligns with trends identified in the NIAH dataset and underscores the challenge of maintaining instruction adherence and coherence over long outputs.
This deviation becomes particularly pronounced when outputs exceed 4,000 tokens, where adherence to instructions significantly diminishes, and further deterioration is observed as outputs approach 16,000 tokens. In contrast, tasks involving shorter outputs, such as those in the NIAH dataset or simpler multi-needle tasks, showed near-perfect performance, highlighting the disparity in model behavior across different sequence lengths.
Potential reasons for this decline include limitations in the self-attention mechanism used in transformers, which may struggle to maintain meaningful context over long sequences. Additionally, models trained with limited long-form data may overfit to shorter patterns, leading to a loss of coherence in extended generations. These findings suggest that architectural changes or improved training strategies may be necessary to overcome these challenges in future iterations of LLMs.
@@
Figure 2: The right side of the figure illustrates the model’s performance on specific instruction tasks at 16K as sequence length increases, whereas the left side depicts performance at 32K. All curves have been smoothed with a Moving Average.
3.4 Three Specific Task Instructions
Figure 3 presents the model’s performance metrics across various task types: single, range, and periodic. The model demonstrated comparable proficiency in both single and range tasks, reflecting its capability to follow direct and straightforward instructions effectively. However, the slight reduction in performance for range tasks suggests that additional complexity, such as processing multiple data points within a defined range, introduces a marginal increase in cognitive load for the model.
The most significant decline in performance was observed in periodic tasks, where the model struggled to interpret instructions that required recurring events, such as ”every four weeks starting from week 10.” These tasks demand a higher degree of reasoning and temporal awareness, which may challenge the model’s capacity to maintain consistency over extended sequences. As a result, outcomes for periodic tasks were considerably poorer compared to single and range tasks, which have clearer and more well-defined parameters. The model’s performance hierarchy can generally be summarized as $single>range>periodic$, highlighting the increased difficulty associated with periodic tasks. This trend underscores the need for future improvements in long-context models, particularly in handling more complex, time-based instructions.
3.5 Comparison with Long-Context Input
@@
We examine the relationship between a model’s ability to handle long inputs and its performance on long outputs. Specifically, we investigate whether a model’s capacity to manage long-range inputs corresponds to improved performance on long-range outputs. For this analysis, we use the RULER dataset, a synthetic benchmark designed to flexibly configure sequence length and task complexity, making it ideal for comprehensive evaluations of long-context LLMs. We compare the models’ performance on sequences of the same length, as shown in Figure 3, which indicates a significant performance gap between input handling and output performance. At 16K tokens, the Pearson correlation coefficient is 0.51, while at 32K tokens, it increases to 0.66, suggesting that there is some overlap in the skills required for managing long inputs and generating long outputs, but these tasks are not entirely equivalent.
Handling long inputs primarily requires the model to retain and process existing information, while generating long outputs demands more complex reasoning, memory retention, and coherence management over extended sequences. Thus, models that excel in long-input retrieval may still struggle with long-form generation, particularly in tasks requiring strict instruction adherence over time. This distinction highlights the need for models to be optimized for both input handling and output generation to achieve consistent performance in long-context tasks.
(a) Performance Comparison on three tasks settings
@@
Figure 3: The left fig displays the models’ performance on three different task settings, with the red line representing the average for each category. The right fig shows the performance and correlation of the Ruler and LongGenBench  at the same length settings.
4 analysis and Limitations
Richness of Content.
Despite efforts to design sub-scenarios that enhance task diversity and richness, the model’s outputs tend to converge as output volume increases. This results in a homogenization of recorded events, even when differences in time and location should introduce variety. Such convergence not only degrades overall performance but also diminishes the diversity of the generated content, leading to repetitive and predictable outputs. In our experiments, approximately 45% of long outputs exhibited significant repetition, even when the model was given varied time or location prompts. Adjusting parameters like repetition_penalty during inference has shown limited success in mitigating this issue, highlighting the need for more advanced techniques to maintain content richness over long sequences.
Rationality of Content.
While our current research focuses primarily on evaluating instruction-following capabilities, a more comprehensive analysis of content rationality and coherence is needed. For example, when tasked with generating a diary, the model should ensure that all recorded activities align with the specified careers. However, in many instances, this logical consistency is lacking. Additionally, temperature records in virtual diary entries often fail to reflect realistic temporal changes. For instance, in a San Francisco’s diary task, we would expect temperatures to vary from cooler (0-10 degrees Celsius) at the beginning of the year to warmer (20-30 degrees Celsius) by mid-year. Yet, the model consistently generates warmer temperatures throughout, even into December. These issues may arise due to the model’s limited exposure to temporally varied datasets, particularly in diary or climate-related contexts. Future work could address this by incorporating more domain-specific and temporally annotated data during fine-tuning.
Instruction Data.
A significant performance discrepancy between models’ abilities to handle long-range inputs (such as Ruler (Hsieh et al., 2024)) and their long-form output generation can likely be attributed to the length distribution of instruction-tuning data. Most instruction-tuning datasets are brief, typically under 200 tokens, and lack the extended instructional content necessary for generating longer outputs. This gap suggests that organizing or synthesizing instruction-tuning data with longer, more comprehensive examples could be a valuable direction for future research. Potential solutions include applying transfer learning techniques from models trained on long-form datasets or using data augmentation methods to synthesize longer instructional content from existing short-form data.
Generalizability.
LongGenBench effectively evaluates instruction-following in creative and technical tasks but may not fully capture the creativity and specialized knowledge required for abstract reasoning or unconstrained storytelling. Future versions could include open-ended tasks like creative fiction writing and legal document drafting, which demand intricate narratives and precision. Expanding in this direction would enhance the benchmark’s versatility while providing deeper insights into LLMs’ capabilities. However, LongGenBench’s current focus on instruction adherence offers a strong foundation for evaluating practical, instruction-driven long-form text generation.
@@ 5 Related Work
Instruction Following.
Recent advances in instruction tuning models (Ouyang et al., 2022; Rafailov et al., 2024; OpenAI, 2022; Taori et al., 2023; Chiang et al., 2023) have underscored the need for scalable evaluation methods. LLMs have been used as evaluators, showing better alignment with human judgments than traditional metrics like BLEU (Papineni et al., 2002). However, LLM evaluations suffer from biases, such as sensitivity to presentation order and preference for verbose outputs (Wang et al., 2024; Pezeshkpour & Hruschka, 2023; Zheng et al., 2023). To mitigate these biases, meta-evaluation benchmarks like FairEval, MT-Bench, and LLMEval2 (Wang et al., 2024; Zheng et al., 2023; Zhang et al., 2023) have been proposed. While recent studies have focused on improving LLM evaluations with diverse strategies (Zheng et al., 2023; Li et al., 2023; Zhang et al., 2023; Chan et al., 2023), they typically do not address longer context lengths.
@@
Existing benchmarks focus on models handling long inputs. For instance, ZeroSCROLLS (Shaham et al., 2023) and LongBench (Bai et al., 2023) tackle tasks like long-document QA and query-based summarization. Synthetic benchmarks, like NeedleBench (Li et al., 2024) and Ruler (Hsieh et al., 2024), offer better control over variables such as sequence length and complexity. NeedleBench introduces the Ancestral Trace Challenge (ATC), while Ruler evaluates models across tasks like NIAH and multi-hop tracing. However, these benchmarks largely focus on input comprehension and do not assess long-form text generation, which is the primary focus of LongGenBench.
@@
Research in long-form generation spans applications like story generation (Fan et al., 2019c; Xu et al., 2020), paragraph completion (Kang & Hovy, 2020), sustained conversation (Xu et al., 2022), and comprehensive QA (Fan et al., 2019a; Dasigi et al., 2021; Stelmakh et al., 2022; Lee et al., 2023). However, existing models and evaluation methods (Liu et al., 2023; Chiang & Lee, 2023; Liu et al., 2024; Bai et al., 2024) face challenges in maintaining quality over long outputs, often being limited by shorter text lengths (typically under 2000 tokens) (Shen et al., 2023). Recent work (Tan et al., 2024) seeks to improve evaluation criteria, but the gap between model capabilities and benchmark text lengths remains. In contrast, LongGenBench evaluates models on their ability to handle much longer sequences, with tasks requiring adherence to instructions over extended outputs (16K+ tokens).
@@ 6 Conclusion
We introduced LongGenBench, a synthetic benchmark that evaluates long-form generation capabilities of language models by testing their ability to follow instructions over extended sequences. In evaluating nine advanced models with context sizes ranging from 32K to 128K tokens, we observed significant performance degradation compared to benchmarks like ”Ruler” with common failure modes including premature task termination, incomplete responses, disregard for instructions, and repetitive content generation. These results highlight key challenges for current models in handling long-form tasks and underscore the need for advancements in model architecture and training data to improve coherence, instruction adherence, and content diversity over extended outputs.
Appendix A PROMPT TEMPLATES FOR FOUR TASK SCENARIOS
Below are the example templates for the four task scenarios: Diary Writing, Menu Design, Skyscraper Design, and Urban Planning. In the Diary Writing template, professions and names are customizable variables, allowing for flexibility in the generated content.
Appendix B evaluation prompts
We provide the prompts used for evaluation and the function $\text{eval}(A_{i},T_{i})$ in Algorithm 1, where Corresponds to the output part of Specific Task Instruction represents $A_{i}$, and Specific Task Instruction represents $T_{i}$. We present both a ”yes” example and a ”no” example for the few-shot evaluation.
Evaluation Prompts Example 1 XXXX Answer: yes Example 2 XXXX Answer: no Context: ”Corresponds to the output part of Specific Task Instruction” Instructions: Does this context include Specific Task Instruction? Please answer with ’yes’ or ’no’ only.”
Appendix C Models
In this benchmark, we evaluated ten LLMs, including both open-source and closed-source models. These models vary in parameter size and context length capabilities, which are crucial factors in their performance on long-form text generation tasks. The key details for each model are outlined in Table 4. These include closed-source models like GPT-4o-mini and GPT-4o, which support a context length of 128K tokens and serve as state-of-the-art baselines for long-context handling. Open-source models, such as Llama3.1-8B and Llama3.1-70B, offer similar context lengths and represent the latest in large-scale, open-access LLMs. Qwen2-7B and Qwen2-72B, developed by Qwen, also support 128K tokens and handle complex long-text tasks. Additionally, we evaluated Mixture of Experts (MoE) models like Mistral-v0.2 and Mixtral-8x7B, both with context lengths of 32K tokens, focusing on memory efficiency and scalability. FILM-7B, designed for creative and technical tasks, supports 128K tokens and excels in generating detailed, context-rich content. Finally, Longwrite-llama3.1-8B, based on Llama3.1, is optimized for long-form narrative tasks with a context window of 128K tokens. Together, these models offer a diverse representation of advancements in long-context LLMs, showcasing their ability to handle long-form, instruction-driven generation tasks.
@@ Appendix C Models
FILM-7B (An et al., 2024) 7B 128K In2Training/FILM-7B
@@
Appendix D Error Analysis
We showcase two error generated output that highlight the common issues faced by large language models in maintaining coherence and instruction-following over extended sequences. In the Diary for 2018 scenario, while the model initially adheres to the instructions for weeks 6 and 7, it progressively diverges from the given timeline, misplacing events and generating off-topic content by week 15 and beyond. Similarly, in the Skyscraper Design task, the model starts with correct descriptions for specific floors, such as Floor 11 designated as an art gallery, but it fails to maintain consistency with the subsequent instructions, particularly in the range and periodic requirements for floors 32-39 and every 15th floor. These examples illustrate the challenges long-context models face in sustaining adherence to task instructions, particularly as sequence length increases.
Diary for 2018 Noah is a photographer with a passion for chronicling his vibrant life through weekly diary entries. Captures in daily: 1) Family member birthday: husband (yourself) (birthday on March 08), wife (birthday on July 09), child_1 (birthday on November 06). 2) Joining a fitness and wellness retreat in Bali in week 6-7. 3) Join a local running club for weekend marathons every 5 weeks on weekends, starting from week 10. ….. =============================== Output =============================== ….. Week 6 (February 5th - February 11th): Noah had joined a fitness and wellness retreat in Bali in this week. …. ✓ Week 7 (February 12th - February 18th): Continuing in Bali, I spent the week participating in various workshops and seminars. …. ✓ ….. Week 15 (April 9th - April 15th): The family is in full spring mode, with kids spending time outside, exploring, and enjoying the nice weather. ….. ✗ ….. Week 35 (August 27th - September 2nd): The summer heat has subsided, and we’ve been enjoying the milder weather 2013 going on hikes, playing in the park, and exploring our neighborhood. …. ✗ ….. Week 45 (November 5th - November 11th): The kids have been enjoying their school routines, and we’ve been making the most of our evenings together. ….. ✗
Skyscraper Design Construct a skyscraper with 100 floors. Please follow the detailed floor assignments below: 1) Specific floors requirements: Designate Floor 11 for small art gallery use. …. 2) Range floors requirements: Allocate Floors 32 to 39 for corporate headquarters for a major company. 3) Periodic floors requirements: Include a sky garden every 15 floors, starting from Floor 30. ….. =============================== Output =============================== ….. Floor 11: Designated for art gallery use, Floor 11 is a sophisticated and flexible space designed to celebrate visual arts….. ✓ ….. Floor 32: Floor 32 serves dual purposes, housing a renowned photography studio and corporate offices. ….. ✗ ….. Floor 34: Transitioning into a leisure space, Floor 34 hosts a small cinema, providing an exclusive entertainment venue within the skyscraper. ….. ✗ ….. Floor 60: This floor houses a luxury watch and timepiece atelier, celebrating the art of horology and fine craftsmanship. ….. ✗ ….. Floor 90: Floor 90 offers a dynamic e-commerce and digital marketing center focused on online business innovation and consumer engagement strategies. ….. ✗